OpenStudy (anonymous):

Suppose a bag of marbles has 4 green, 2 red, 5 yellow, 1 brown, and 7 blue marbles. What is the probability of picking a green marble?

OpenStudy (anonymous):

Total number of marbles = 4+2+5+1+7 = 19 p(green) = 4/19